OpenStudy (zmudz):

Let \(n\) be a positive integer. Show that the smallest integer greater than \((\sqrt{3} + 1)^{2n}\)is divisible by \(2^{n+1}.\) Hint: Prove that \(\lceil (\sqrt{3}+1)^{2n} \rceil = (\sqrt{3}+1)^{2n} + (\sqrt{3}-1)^{2n}.\)

OpenStudy (amilapsn):

you can use mathematical induction with binomial expansion.

OpenStudy (anonymous):

Id probably say make a recurrence relation

OpenStudy (zzr0ck3r):

A number can't be equal to itself plus a nonzero number...

OpenStudy (anonymous):

Isn't that\[\lceil (\sqrt{3}+1)^{2n} \rceil = (\sqrt{3}+1)^{n} + (\sqrt{3}-1)^{n}.\]

OpenStudy (amilapsn):

\[\exists ! 0.abcd.... ~s.t.\] \[\lceil{(\sqrt{3}+1)^{2n}}\rceil=(\sqrt{3}+1)^{2n}+0.abcd.....\] \[But~ 0<(\sqrt{3}-1)^{2n}< 1\forall n\in \mathbb{Z}^+\] By binomial expansion: \[\small (\sqrt{3}+1)^{2n}=^{2n}C_0 (\sqrt{3})^{2n}+^{2n}C_1 (\sqrt{3})^{2n-1}+\ldots+^{2n}C_{2n-1} (\sqrt{3})^{1}+^{2n}C_{2n} 1 ~~(2n+1~Terms)\] \[\small (\sqrt{3}-1)^{2n}=^{2n}C_0 (\sqrt{3})^{2n}-^{2n}C_1 (\sqrt{3})^{2n-1}+\ldots-^{2n}C_{2n-1} (\sqrt{3})^{1}+^{2n}C_{2n} 1 ~~(2n+1~Terms)\] \[\small \therefore (\sqrt{3}+1)^{2n}+(\sqrt{3}-1)^{2n}=2 \left[^{2n}C_0 (\sqrt{3})^{2n}+^{2n}C_{2} (\sqrt{3})^{2n-2}+\ldots+^{2n}C_{2n-2} (\sqrt{3})^{2}+^{2n}C_{2n} 1\right] \] Since all the terms have even exponents of \(\sqrt{3}\) this will be an integer... \[\therefore \lceil{(\sqrt{3}+1)^{2n}}\rceil=(\sqrt{3}+1)^{2n}+(\sqrt{3}-1)^{2n}\] By mathematical induction we can show that.. \[2^{n+1}|\lceil{(\sqrt{3}+1)^{2n}}\rceil\]

OpenStudy (zzr0ck3r):

suppose \(a=a+b\) where \(b\ne 0\). Then by cancelation \(b=0\) a contradiction. :)

OpenStudy (amilapsn):

@zzr0ck3r you've forgotten the ceiling... :) for example \(\lceil 2.5\rceil =3\)

OpenStudy (zzr0ck3r):

Ahh, sorry on my phone and I thought it was just \([\). lol. I was going nuts...